# Probing loop quantum effects through solar system experiments: observational signatures and parameter constraints

### v2

@@ I Introduction
In this paper, we investigate a LQG-inspired BH spacetime featuring double horizons where quantum gravitational effects parameterized by the dimensionless deformation parameter $\zeta$ Zhang:2024khj . While prior studies have extensively characterized this spacetime’s strong-field phenomenology through quasi-normal mode (QNM) spectrum Konoplya:2024lch ; Skvortsova:2024msa ; Zhang:2024svj , photon rings, shadow morphology Liu:2024soc ; Liu:2024wal ; Xamidov:2025oqx , spinning particle dynamics Du:2024ujg , and accretion disk structures Shu:2024tut ; Chen:2025ifv , etc. The BH solution has also been extended to rotating counterparts Ban:2024qsa and scenarios incorporating vacuum charges and cosmological constants Yang:2025ufs . Although prior studies have constrained this model through strong-field observations Konoplya:2024lch ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv ; Xamidov:2025oqx , systematic parameter constraints from weak-field regimes, particularly solar system experiments, remain conspicuously absent. We therefore leverage high-precision solar system tests of GR to establish the observational bounds on $\zeta$ [^1] During the final stages of this work, we became aware that constraints on the parameter $\zeta$ have been independently investigated in Xamidov:2025oqx using the MESSENGER mission data from Mercury’s relativistic perihelion precession and the pericenter advance of the S2 star orbiting Sgr $\rm{A}^{*}$..
This paper is organized as follows. Section II provides a thorough geometric characterization of the effective LQG-BH spacetime, with particular emphasis on the geodesic motion of particles in its exterior spacetime. In Section III, the classical GR experiments are employed to probe the effects of the quantum gravity effect. Section IV summarizes the key findings and outlines prospective directions for future investigations. Throughout this work, we adopt geometrized Planck units unless otherwise specified. When using experimental data, the International System of Units (SI) is restored for calculations.
@@ II Test Particle Dynamics in an Effective LQG-BH Spacetime
In this section, we begin by providing a concise overview of the effective LQG-BH model, parameterized by a quantum parameter $\zeta$. We then derive the equations of motion (EOMs) for a test particle orbiting the BH.
@@ II.1 An effective LQG-BH spacetime
Here, the LQG-corrected parameter $\zeta$ is defined as
$\zeta=\frac{\sqrt{4\sqrt{3}\pi\gamma^{3}\ell_{\mathrm{P}}^{2}}}{M}\,,$ (3)
where $M$ denotes the BH mass, $\gamma$ is the Barbero-Immirzi (BI) parameter of LQG, and $\ell_{\mathrm{P}}$ represents the Planck length. Crucially, since the BI parameter lacks a first-principles determination in current LQG frameworks and is conventionally treated as a free parameter Meissner:2004ju ; Domagala:2004jt , we consequently adopt $\zeta$ as an effective dimensionless free parameter throughout our analysis. This approach allows systematic investigation of LQG-induced modifications to BH.
@@ III.1 Deflection of light
where the factor of $2$ accounts for symmetric deflection during approach and recession. To evaluate the integral in the above equation, we implement the dimensionless substitution $u=\frac{r_{0}}{r}$. In the weak-field regime characterized by $\epsilon\equiv\frac{M}{r_{0}}\ll 1$, we perform a perturbative expansion of $\bigtriangleup\phi$ to second order in $\epsilon$. This yields the asymptotic expression:
@@
where $\Delta\phi_{\mathrm{GR}}$ represents the standard deflection in GR, with a value of approximately $1.75$ arcsec. We note that the parameter $\epsilon$ has been restored to $\frac{M}{r_{0}}$ in the above expression.
@@ III.3 Precession of perihelia
$u\left(\phi\right)\approx\frac{Mr_{0}}{J^{2}}\left(1+e\cos\phi\right)+\left(\frac{\chi_{1}}{2}+\frac{3\chi_{3}}{8}\right)\phi\sin\phi\approx\frac{Mr_{0}}{J^{2}}\left\{1+e\cos\left[\left(1-\frac{\delta\phi}{2\pi}\right)\phi\right]\right\}\,,$ (42)
@@ IV Conclusion
Table 1: Upper bounds on quantum-gravity-corrected parameter $\zeta$ in the effective LQG black hole spacetime from solar system and galactic center data.
 Experiments/Observations $\zeta$  Datasets
Light deflection 9.126139.126139.12613$\quad$9.12613   VLBI observation of quasars
Shapiro time delay 2.609862.609862.60986$\quad$2.60986   Cassini experiment
2.602.602.60$\quad$2.60  Doppler tracking of Cassini
Perihelion advance 0.01120890.01120890.0112089$\quad$0.0112089  MESSENGER mission
0.1056410.1056410.105641$\quad$0.105641  LAGEOS satellites
0.7348470.7348470.734847$\quad$0.734847   Observation of S2 around Sgr $\rm{A}^{*}$
@@
Theoretical calculations reveal that the inclusion of ${\zeta}^{2}$ terms induces deviations from GR predictions, resulting in a lagged manifestation of relativistic phenomena. As summarized in Table 1, the tightest constraint on $\zeta$ arises from the MESSENGER mission data of the Mercury periasis shift, yielding $0.01120890<\zeta<$ 0.0112089, whereas the second most stringent constraint is produced by the LAGEOS satellites, at the level of $10^{-1}$. To further constrain $\zeta$, we employ the strong gravitational field observations of the S2 star orbit around Sgr $\rm{A}^{*}$, deriving an upper bound $\zeta\lesssim 10^{-1}$, yielding a tighter constraint than those obtained from EHT data of BH shadow radius Konoplya:2024lch ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv . Additionally, we estimate the theoretical value of the rescaled parameter $\zeta$ (Eq. (3)) to depend on the solar mass by fixing the BI parameter $\gamma=0.2375$, which gives $\zeta\sim 10^{-39}$. While current solar system experiments lack the resolution to detect LQG signatures, next-generation GW detectors (e.g., LISA LISA:2022yao , Einstein Telescope ET:2019dnz ) may probe such signatures and impose stronger constraints. Furthermore, we anticipate ongoing missions like Gaia Gaia:2016zol and BepiColombo Will:2018mcj , along with proposed space-based projects such as LATOR LATOR:2005hia and BEACON Turyshev:2008rh , will progressively improve the precision of LQG-corrected parameters constraints.

### v3

@@ I Introduction
[…]
In this paper, we will probe quantum gravity effects through solar system experiments in a LQG-inspired BH spacetime featuring double horizons, parameterized by dimensionless deformation $\zeta$ Zhang:2024khj . The paper is organized as follows. Section II provides a detailed geometric characterization of the effective LQG-BH spacetime, with particular emphasis on the geodesic motion of particles in its exterior spacetime. In Section III, the classical GR experiments are employed to probe the effects of the quantum gravity effect. Section IV summarizes the key findings and describes prospective directions for future investigations. Throughout this work, we adopt geometrized Planck units unless otherwise specified. When using experimental data, the International System of Units (SI) is restored for calculations.
@@ II Test Particle Dynamics in an Effective LQG-BH Spacetime
In this section, we first outline the effective LQG-BH spacetime Zhang:2024khj , then derive the equations of motion (EOMs) for test particles orbiting this quantum-corrected geometry.
@@ II.1 An effective LQG-BH spacetime
The dimensionless parameter $\zeta$ encodes quantum gravitational corrections and is fundamentally defined as:
$\zeta=\frac{\gamma\sqrt{A}}{M}\,$ (3)
where $M$ is the BH mass, $\gamma$ denotes the Barbero-Immirzi (BI) parameter of LQG, and $A$ represents the minimal area gap from LQG’s holonomy quantization. Specifically, $A=4\sqrt{3}\pi\gamma\ell_{\mathrm{P}}^{2}$ corresponds to the smallest non-zero eigenvalue of LQG’s area operator Ashtekar:2006wn , with $\ell_{\mathrm{P}}$ being the Planck length. This parameter $\zeta$ determines the onset of quantum gravity effects, where the classical singularity ($A\to 0$) is recovered as $\zeta\to 0$. Since the BI parameter $\gamma$ currently lacks first-principles determination in current LQG frameworks Meissner:2004ju ; Domagala:2004jt , we treat $\zeta$ as an effective dimensionless free parameter in our phenomenological approach. This approach allows systematic investigation of LQG-induced modifications to BH.
@@ III.1 Deflection of light
where the factor of 222 accounts for symmetric deflection during approach and recession. To evaluate the integral in the above equation, we implement the dimensionless substitution $u=\frac{r_{0}}{r}$. In the weak-field regime characterized by $\epsilon\equiv\frac{M}{r_{0}}\ll 1$, we perform a perturbative expansion of $\bigtriangleup\phi$ to second order in $\epsilon$. This yields the asymptotic expression:
@@
where $\Delta\phi_{\mathrm{GR}}$ represents the standard deflection in GR, with a value of approximately 1.751.751.75 arcsec. We note that the parameter $\epsilon$ has been restored to $\frac{M}{r_{0}}$ in the above expression.
@@ III.3 Precession of perihelia
$u\left(\phi\right)\approx\frac{Mr_{0}}{J^{2}}\left(1+e\cos\phi\right)+\left(\frac{\mathcal{P}_{1}}{2}+\frac{3\mathcal{P}_{3}}{8}\right)\phi\sin\phi\approx\frac{Mr_{0}}{J^{2}}\left\{1+e\cos\left[\left(1-\frac{\delta\phi}{2\pi}\right)\phi\right]\right\}\,,$ (42)
@@
Before concluding this section, we note that during the final stages of this work, an independent study by Xamidov:2025oqx constrained the parameter $\zeta$. Their analysis used the estimated result of Mercury’s relativistic perihelion precession given by 42.980±0.00242.980$\pm$ 0.00242.980 ± 0.002 arcsec per century Benczik:2002tt , and the GRAVITY data of the pericenter advance of the S2 star orbiting Sgr $\rm{A}^{*}$, reporting $\zeta\leq 0.01869$ (Mercury) and $\zeta\leq 0.73528$ (S2 star), both consistent with our constraints. Additionally, by applying quasiperiodic oscillation (QPO) data from four astrophysical sources (GRO J1655-40, XTE J1550-564, GRS 1915+105, H1743-322) and Markov Chain Monte Carlo (MCMC) analysis in strong-gravity regimes, they derived $\zeta\leq 2.086$ — a result significantly tighter than previous bounds from BH shadow observations.
In comparison, our work incorporates light deflection measurements from VLBI observations of quasars and Shapiro time-delay data from the Cassini experiment (including Doppler tracking), providing additional independent bounds. This alignment of independently obtained bounds underscores the robustness of current limits on the LQG parameter $\zeta$ and demonstrates consistency across distinct methodologies.
@@ IV Conclusion
Table 1: Summary of observational constraints on the LQG-corrected parameter $\zeta$. Datasets are drawn from solar system and galactic center observations, with corresponding relative uncertainties.
 Experiments/Observations  Relative uncertainties  $\zeta$  Datasets
[…]
/ 2.602.60$\quad$2.60  Doppler tracking of Cassini
[…]
@@
Theoretical calculations reveal that the inclusion of ${\zeta}^{2}$ terms induces deviations from GR predictions, resulting in a lagged manifestation of relativistic phenomena. As summarized in Table 1, the tightest constraint on $\zeta$ arises from the MESSENGER mission data of the Mercury periasis shift, yielding $0<\zeta<0.0112089$, whereas the second most stringent constraint is produced by the LAGEOS satellites, at the level of $10^{-1}$. To further constrain $\zeta$, we employ the strong gravitational field observations of the S2 star orbit around Sgr $\rm{A}^{*}$, deriving an upper bound $\zeta\lesssim 10^{-1}$, yielding a tighter constraint than those obtained from EHT data of BH shadow radius Konoplya:2024lch ; Heidari:2024bkm ; Wang:2024iwt ; Shu:2024tut ; Chen:2025ifv .
It is valuable to contextualize theses observational bounds with theoretical expectations. Fixing the BI parameter at $\gamma=0.2375$ Domagala:2004jt , Eq. (3) gives the characteristic scaling:
$\zeta_{\mathrm{theo}}^{\odot}\sim\frac{\gamma\thinspace\mathcal{O}\left(A^{\frac{1}{2}}\right)}{{M}_{\odot}}\approx 10^{-39}\,,$ (53)
where we substitute $A=4\sqrt{3}\pi\gamma\ell_{P}^{2}$ with $\ell_{P}\sim 10^{-35}$ m and solar mass ${M}_{\odot}\sim 10^{30}$ kg. This reveals the critical mass dependence of quantum gravity effects:
  • Theoretical $\zeta\propto 1/M$ explains why solar-system constraints ($\zeta<0.01$) remain 37$\sim$ 37∼ 37 orders above $\zeta_{\mathrm{theo}}^{\odot}$.
  • Quantum gravity effects become significant when $M\lesssim\gamma\sqrt{A}/\zeta_{\mathrm{obs}}\sim(10^{-6}-10^{-8})$ kg (Planck-mass scales).
While current solar-system tests cannot probe fundamental $\zeta$ values, our constraints establish empirical upper bounds for phenomenological LQG models. Future work will prioritize strong-field regimes, particularly GW signatures from BH mergers.
Although current solar system experiments lack the sensitivity to detect LQG signatures, upcoming space-based missions will progressively tighten constraints on quantum-corrected parameters. For instance, the optical interferometry-based Gaia orbital telescope Gaia:2016zol , with its superb light deflection measurements, is projected to achieve the PPN parameter $\gamma$ with an accuracy in the range of $10^{-5}\sim 10^{-7}$. The Mercury Orbiter Radio-science Experiment (MORE) of BepiColombo will exploit solar conjunctions to measure radio gravitational delays, with predicted $\gamma$ constraints reaching $10^{-6}$ level Will:2018mcj ; Genova:2021rlo . Furthermore, complementary efforts by Jupiter orbiting missions such as Juno and JUICE, will further refine $\gamma$ estimates via spacecraft trajectory tracking, with the latter targeting precision at the $10^{-7}$ level DeMarchi:2019lei . Synthesizing these multi-mission datasets could improve current $\zeta$ bounds by 1–2 orders of magnitude. Meanwhile, next-generation GW detectors (e.g., LISA LISA:2022yao , Einstein Telescope ET:2019dnz ) may directly probe such signatures and impose even stronger constraints.